### v3

# Sun-Shine: A Foundation Large Language Model for Tibetan Culture and Heritage

Abstract
Tibetan, a minority language in China, features a highly intricate grammatical structure, characterized by four verb tenses and a tense system with frequent irregularities, contributing to its extensive inflectional diversity. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in many domains. Despite the success in other fields, current LLMs often fall short in catering to the needs of domain experts like Tibetans, and the potential of LLMs for Tibetan culture is under-explored. The intrinsic reasons are the immense and intricate nature of Tibetan culture as well as the necessity for higher granularity and richness in knowledge. Simultaneously, the complexity and uniqueness of its grammatical structure, coupled with its status as a minority ethnic language, contribute to data scarcity, which remains a fundamental challenge. To alleviate these issues, we introduce Llama-Sunshine (Sun-Shine), the first large language model for Tibetan culture, which is expert in various Tibetan language processing tasks. Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features. We also propose TIB-STC, a comprehensive benchmark comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale benchmark for Tibetan culture. Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis. Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities.
Index Terms:
Large Language Model, Foundation Model, Tibetan Language Benchmark, Tibetan Language Processing1 Introduction
Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT [1], Gemini [12], Qwen-2.5 [34, 33], Llama [2, 20] and DeepSeek [37, 38]. Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges [4]. Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA [2] and DeepSeek [37, 38], which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
Despite their many strengths, LLMs are not tailored to specific fields. Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain [3, 5]. Also, there are some biases against LLMs [6, 7, 8]. LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage [4]. For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments [9]. Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability. Overall, data scarity is the key issue for these areas. Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem [11]. In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language [10]. Assembling a benchmark about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort. Only then can we develop a more specialized and unbiased LLM model on this basis.
To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus benchmark designed to address the issue of data scarcity in Tibetan language research. This benchmark spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan benchmark in the world. On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture. the first LLM specifically tailored for Tibetan language processing. During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 [38] totally, showing a higher level of knowledge expertise in Tibetan Culture. Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture. Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and benchmark for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
In summary, our contributions can be summarized as follows:
We present Sun-Shine, the first LLM for Tibetan culture, leveraging Chinese-to-Tibetan translation and manual annotation to tackle linguistic complexities and enhance training accuracy, paving the way for NLP in low-resource languages.
We introduce TIB-STC, the first large-scale Tibetan benchmark, addressing data scarcity in Tibetan language research. Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our benchmark, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
2 Related Work
2.1 Tibetan Language Processing
LLMs for rare or low-resource languages are relatively uncommon [13]. The absence of corpus data is the primary and most critical issue [9, 11] like Tibetan. However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently [16]. As a result, digital Tibetan language resources are exceptionally scarce. This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing [11, 10, 26, 19].
For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Fig. 1.
As shown in Fig. 2, the native Tibetan syllable structure, includes seven kinds of letters. There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel. Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data. To solve this issue, we propose the first large-scale Tibetan benchmark, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
2.2 Large Language Model
LLMs, including GPT [4], Gemini [12], and ChatGLM [6, 7], are highly parameterized models with billions of parameters, pre-trained on massive natural language benchmarks [14]. Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities. By applying task-specific instructions and fine-tuning techniques [15, 17, 18], LLMs can be optimized for specialized fields, enhancing their performance and domain expertise [5].
Taking into account the unique characteristics of Tibetan culture and in order to make life more convenient for the Tibetan people, we introduce the first LLM for Tibetan, Sun-Shine which is based on TIB-STC and cover a wide range of subjects, such as medicine, law, sports, daily life and basic sciences, providing adequate support can significantly improve the daily lives of people in Tibetan areas.
3 TIB-STC and Sun-Shine
3.1 Benchmark: TIB-STC
TIB-STC, is the first comprehensive and detailed Tibetan corpus benchmark for LLM, consisting of five categories of data as shown in the benchmark branch of Table I.
| Benchmark | Instance | Tokens | Storage Size |
| Corpus | 122,756,184 | 11B | 18GB |
| Alpaca-Ti | 42,676 | 76M | 109MB |
| Safety-Prompts-Ti | 38,156 | 62M | 116MB |
| CValues-Ti | 2,000 | 6M | 9MB |
| hh-rlhf-Ti | 47,411 | 98M | 154MB |
Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement. For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability. Pre-Training involves the largest benchmark in terms of instances, tokens, and storage, as it focuses on general language model training. In the stage of Supervised Fine-Tuning (SFT), benchmarks are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety). Direct Preference Optimization (DPO) uses the smallest benchmarks but focuses on high-quality preference data for alignment and optimization.
For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
3.1.1 Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances.
Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation. Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base. As shown in Fig. 3, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
3.1.2 Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from Alpaca [31, 32] project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table II.
| Domain | Category |
| Daily Conversation | Small Talk and Q&A |
| Education | Multidisciplinary Problem and Solving |
| Coding | Code Generation and Debugging |
| Creative Writing | Poetry, Story |
| Practical Tasks | Summary, Translation, Text Proofreading |
| Scientific Reasoning | Logical Reasoning and Calculation |
| Areas of Expertise | Medicine, Law, Finance, etc. |
| Entertainment | Riddles, Jokes, Game Suggestions. |
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Fig. 4.
3.1.3 Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table III, and the data size is 38,156. LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts [22, 23, 24].
| Instruction Attacks | 15k |
| Instruction Attacks consists of Goal Hijacking | 2.5K |
| Prompt Leaking | 2.5K |
| RolePlay Instruction | 2.5K |
| Unsafe Instruction Topic | 2.5K |
| Inquiry with Unsafe Opinion | 2.5K |
| Reverse Exposure | 2.5K |
| Typical Safety Scenarios | 23K |
| Typical Safety Scenarios consists of Insult | 3.6K |
| Unfairness and Discrimination | 3.8K |
| Crimes and Illegal Activities | 3.6K |
| Physical Harm | 1.7K |
| Mental Health | 4K |
| Privacy and Property | 1.3K |
| Ethics and Morality | 1.3K |
| Others | 4.7K |
Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Fig. 5.
3.1.4 CValues-Ti
CValues-Ti is the benchmark, used to improve the security of Tibetan language models. Considering work [25, 27, 28], we propose this branch of a small benchmark because LLM can pose risks or has negative social impacts. For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work [28], We streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Fig. 6.
3.1.5 hh-rlhf-Ti
hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF [29, 30], but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Fig. 8.
3.2 Sun-Shine
Sun-Shine is built on the LLaMA-3 [2, 20] architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model [20]. Additionally, we integrate NLP model translation with human annotation [21] as supplementary data to enhance Sun-Shine’s accuracy.
As illustrated in Fig. 7, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). For training details, Table IV has shown parameters’ setting. Notably, we adopted LoRA-based parameter-efficient fine-tuning [15] for both SFT and DPO, which significantly reduced memory usage without sacrificing performance. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
| Parameter | Training Stage | ||
| Pre-Training | SFT | DPO | |
| Precision | bfloat16 | bfloat16 | bfloat16 |
| Optimizer | AdamW | AdamW | AdamW |
| Learning Rate | |||
| LR Scheduler | CosineAnnealing | CosineAnnealing | CosineAnnealing |
| Warmup Ratio | 0.1 | 0.1 | 0.1 |
| Gradient Accumulation | 4 | 4 | 8 |
| Epochs | 3 | 3 | 3 |
| Batch Size (tokens) | 4096 | 4096 | 4096 |
| Max Seq Length | 4096 | 4096 | 4096 |
| LoRA | No | Yes (rank=32) | Yes (rank=32) |
| Weight Decay | 0.1 | 0.1 | 0.1 |
| Dropout | 0.1 | 0.1 | 0.1 |
| pref_beta | – | – | 0.1 |
| pref_loss | – | – | Sigmoid |
3.2.1 Pre-Training
During the Pre-training stage, the Sun-Shine is pre-trained on Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction. And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text. For details,
3.2.2 Supervised Fine-Tuning
Following pretraining, the model undergoes supervised fine-tuning [18] with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
3.2.3 Direct Preference Optimization
In the stage of DPO [17, 22, 23], the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
4 Evaluation: TLUE Benchmark
Unfortunately, existing benchmarks like CMMLU [35] and Safetybench [36, 24] are available only in Chinese or English, with no purely Tibetan versions. To address this issue, we used TLUE [42], the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench. Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
4.1 Ti-MMLU
CMMLU [35] is a well-established Chinese benchmark that covers multiple disciplines, including natural sciences, social sciences, engineering, and the humanities. To adapt this benchmark for Tibetan, TLUE [42] leveraged an extended version, Ti-MMLU, where the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance. It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark includes 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation framework follows the original CMMLU metrics, including Response Rate, Accuracy, and Conditional Accuracy. Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
4.2 Ti-Safetybench
SafetyBench [36, 24] is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, and Privacy and Property. Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples. To evaluate safety-related capabilities for Tibetan, TLUE [42] employed Ti-Safetybench, a Tibetan adaptation of SafetyBench. The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards. The evaluation methodology incorporates two scoring methods: Direct Answer Calculation and Concern All Answer Calculation. DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations. Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
Note: as shown in Table V, in order to make the paper more aesthetically pleasing and easier for researchers to read, some specialized terms are abbreviated.
| Category | Full Name | Abbreviation |
| STEM | STEM | |
| Humanities | Human | |
| Ti-MMLU | Social Sciences | Social |
| Other | Other | |
| China Specific | China | |
| Offensiveness | OFF | |
| Unfairness and Bias | UB | |
| Physical Health | PH | |
| Ti-Safetybench | Mental Health | MH |
| Illegal Activities | IA | |
| Ethics and Morality | EM | |
| Privacy and Property | PP | |
| Response Rate | RR | |
| Accuracy | ACC | |
| Evaluation Metrics | Conditional Accuracy | CA |
| Direct Answer Calculation | DA | |
| Concern All Answer Calculation | CAA |
5 Experimental Result
We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) [4], Llama-3.18B [20, 2], Qwen-2.5 (72B) [34] and DeepSeek-R1 [37, 38]. Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis. As shown in Table VI, regarding open-source model execution, the Qwen-2.5 and DeepSeek-r1 were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.18B was accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
| LLM | Temperature | Top_p | Stream |
| GPT-3.5-turbo | 1.0 | 1.0 | False |
| GPT-4o | 1.0 | 1.0 | False |
| GPT-o1-mini | 1.0 | 1.0 | True |
| DeepSeek-R1 | 1.0 | None | Rtue |
| LlaMA-3.18B | 0.6 | 0.9 | False |
| Qwen-2.5 | 0.7 | 0.8 | False |
Also, as shown in Table VII, LLM and some indicators of downstream tasks, such as BLUE [40] and ROUGH [41], used to evaluate translation tasks, also uses abbreviations in some tables for easier reading. For the experimental metrics, the best metric is bolded and the suboptimal metric is underlined.
| Category | Full Name | Abbreviation |
| LLM | GPT-3.5-turbo | G-35T |
| GPT-4o | G-4o | |
| GPT-o1-mini | Go1-m | |
| DeepSeek-R1 | DK-r1 | |
| LlaMA-3.18B | L-318B | |
| Qwen-2.5 | Q-2.5 | |
| Sun-Shine | S-S | |
| Evaluation Metrics | BLEU-4 | B-4 |
| ROUGE-1 | R-1 | |
| ROUGE-2 | R-2 | |
| ROUGE-l | R-l |
5.1 Insights from Performance Results
Ti-MMLU: as shown in Fig. 9, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs.
For Response Rate, as shown in Table VIII, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1. Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training benchmark.
| Category | Response Rate | ||||||
| S-S | G-4o | L-318 | G-35T | Q-2.5 | Go1-m | DK-r1 | |
| ancient_chinese | 0.5366 | 0.4695 | 0.1402 | 0.0732 | 0.2500 | 0.2317 | 0.3232 |
| chinese_civil_service_exam | 0.4938 | 0.4062 | 0.1688 | 0.0438 | 0.2312 | 0.1000 | 0.3812 |
| chinese_food_culture | 0.5588 | 0.5147 | 0.1250 | 0.0809 | 0.2279 | 0.1765 | 0.4191 |
| chinese_foreign_policy | 0.5206 | 0.4981 | 0.3458 | 0.0093 | 0.2804 | 0.1589 | 0.4579 |
| chinese_teacher_qualification | 0.5972 | 0.6089 | 0.2682 | 0.0726 | 0.2458 | 0.3073 | 0.4078 |
| elementary_chinese | 0.5238 | 0.4167 | 0.1667 | 0.0556 | 0.2817 | 0.1667 | 0.4683 |
| ethnology | 0.5481 | 0.5037 | 0.1852 | 0.0963 | 0.3259 | 0.2148 | 0.6148 |
| high_school_politics | 0.5105 | 0.4955 | 0.1888 | 0.0979 | 0.2098 | 0.1748 | 0.4685 |
| modern_chinese | 0.5431 | 0.5241 | 0.1724 | 0.0862 | 0.2069 | 0.2328 | 0.3707 |
| chinese_literature | 0.6127 | 0.5412 | 0.1618 | 0.0833 | 0.2892 | 0.2500 | 0.5000 |
| chinese_history | 0.5635 | 0.4489 | 0.1672 | 0.0526 | 0.2632 | 0.1734 | 0.3839 |
| chinese_driving_rule | 0.5198 | 0.4420 | 0.1985 | 0.0687 | 0.2519 | 0.2595 | 0.3817 |
| construction_project_management | 0.6619 | 0.4388 | 0.1942 | 0.0647 | 0.2662 | 0.2230 | 0.3453 |
| elementary_commonsense | 0.5404 | 0.5354 | 0.1414 | 0.0758 | 0.2828 | 0.1465 | 0.4848 |
| traditional_chinese_medicine | 0.5459 | 0.5405 | 0.2378 | 0.0703 | 0.2324 | 0.2595 | 0.3027 |
| business_ethics | 0.5359 | 0.5167 | 0.2057 | 0.0574 | 0.2919 | 0.2249 | 0.3780 |
| economics | 0.5283 | 0.4465 | 0.2767 | 0.0440 | 0.2516 | 0.1824 | 0.3396 |
| education | 0.5460 | 0.5215 | 0.1840 | 0.0920 | 0.2761 | 0.3129 | 0.4049 |
| college_education | 0.5981 | 0.5234 | 0.2336 | 0.1589 | 0.4299 | 0.2617 | 0.4673 |
| journalism | 0.5947 | 0.5447 | 0.1744 | 0.0407 | 0.3256 | 0.2209 | 0.4826 |
| marketing | 0.6556 | 0.5944 | 0.2278 | 0.0889 | 0.2667 | 0.3500 | 0.2389 |
| professional_accounting | 0.6114 | 0.5829 | 0.1829 | 0.0800 | 0.1943 | 0.2400 | 0.3714 |
| professional_psychology | 0.5645 | 0.5303 | 0.2198 | 0.0819 | 0.1940 | 0.2586 | 0.3578 |
| public_relations | 0.5920 | 0.5172 | 0.1437 | 0.0862 | 0.2299 | 0.2931 | 0.4598 |
| security_study | 0.5704 | 0.4963 | 0.1630 | 0.1185 | 0.4074 | 0.3037 | 0.4000 |
| high_school_geography | 0.5678 | 0.4831 | 0.2712 | 0.0678 | 0.3305 | 0.1780 | 0.4237 |
| management | 0.5905 | 0.5857 | 0.2524 | 0.0857 | 0.2429 | 0.3190 | 0.3810 |
| sociology | 0.5442 | 0.5177 | 0.1991 | 0.0442 | 0.2832 | 0.3319 | 0.3805 |
| marxist_theory | 0.6085 | 0.5556 | 0.2804 | 0.0529 | 0.3122 | 0.2804 | 0.4497 |
| college_law | 0.4648 | 0.4519 | 0.2315 | 0.0833 | 0.2870 | 0.2500 | 0.3981 |
| global_facts | 0.5235 | 0.4930 | 0.2282 | 0.0738 | 0.3423 | 0.2349 | 0.4027 |
| international_law | 0.5568 | 0.5081 | 0.2270 | 0.0595 | 0.3351 | 0.2541 | 0.0919 |
| jurisprudence | 0.5377 | 0.5109 | 0.3187 | 0.0681 | 0.3358 | 0.2409 | 0.3869 |
| world_religions | 0.5000 | 0.4938 | 0.1125 | 0.0938 | 0.2625 | 0.1812 | 0.4688 |
| logical | 0.5366 | 0.4715 | 0.1463 | 0.0976 | 0.2195 | 0.2195 | 0.2846 |
| professional_law | 0.5071 | 0.4882 | 0.2654 | 0.0474 | 0.2180 | 0.2180 | 0.3033 |
| philosophy | 0.5524 | 0.4667 | 0.2476 | 0.0952 | 0.2190 | 0.2286 | 0.0476 |
| world_history | 0.5342 | 0.4907 | 0.2174 | 0.0683 | 0.3354 | 0.2298 | 0.4720 |
| arts | 0.0750 | 0.5437 | 0.1500 | 0.1000 | 0.2375 | 0.2125 | 0.4500 |
| electrical_engineering | 0.5698 | 0.5349 | 0.2035 | 0.1337 | 0.2965 | 0.2384 | 0.3663 |
| actuarial_science | 0.3692 | 0.1887 | 0.1698 | 0.1792 | 0.0642 | 0.1038 | 0.5000 |
| engineering_hydrology | 0.5755 | 0.5377 | 0.2170 | 0.0660 | 0.3113 | 0.3208 | 0.5189 |
| genetics | 0.5455 | 0.4830 | 0.1932 | 0.0795 | 0.1648 | 0.1761 | 0.2216 |
| astronomy | 0.5758 | 0.4606 | 0.2424 | 0.0727 | 0.1455 | 0.2182 | 0.3879 |
| college_mathematics | 0.5048 | 0.3238 | 0.2095 | 0.1429 | 0.1619 | 0.1524 | 0.1429 |
| college_medical_statistics | 0.6038 | 0.5094 | 0.1698 | 0.0566 | 0.1792 | 0.1887 | 0.0660 |
| virology | 0.5680 | 0.5089 | 0.1479 | 0.0414 | 0.1893 | 0.2899 | 0.2071 |
| computer_science | 0.5686 | 0.4216 | 0.1961 | 0.0882 | 0.1225 | 0.1814 | 0.2696 |
| conceptual_physics | 0.5986 | 0.5918 | 0.3878 | 0.0340 | 0.1565 | 0.2925 | 0.2109 |
| anatomy | 0.6622 | 0.5203 | 0.2027 | 0.0676 | 0.1689 | 0.2635 | 0.2905 |
| machine_learning | 0.5656 | 0.4754 | 0.1967 | 0.0574 | 0.1639 | 0.1393 | 0.2951 |
| high_school_physics | 0.6364 | 0.5455 | 0.1909 | 0.0727 | 0.2455 | 0.2198 | 0.3415 |
| high_school_mathematics | 0.5915 | 0.2317 | 0.1646 | 0.2012 | 0.1341 | 0.1098 | 0.3415 |
| high_school_chemistry | 0.5758 | 0.3409 | 0.2955 | 0.0682 | 0.1591 | 0.1288 | 0.0985 |
| high_school_biology | 0.6746 | 0.4852 | 0.2604 | 0.0828 | 0.1657 | 0.2249 | 0.0947 |
| elementary_mathematics | 0.5996 | 0.3870 | 0.1913 | 0.1565 | 0.1696 | 0.1478 | 0.5478 |
| legal_and_moral_basis | 0.5748 | 0.5561 | 0.2523 | 0.0888 | 0.3131 | 0.2523 | 0.5514 |
| computer_security | 0.5789 | 0.5088 | 0.2398 | 0.1053 | 0.2105 | 0.2339 | 0.2164 |
| food_science | 0.4895 | 0.4705 | 0.1469 | 0.0909 | 0.2587 | 0.1818 | 0.3077 |
| college_medicine | 0.4982 | 0.4908 | 0.2674 | 0.0366 | 0.1722 | 0.1538 | 0.1758 |
| clinical_knowledge | 0.5738 | 0.4557 | 0.1899 | 0.1013 | 0.2447 | 0.2616 | 0.1308 |
| professional_medicine | 0.5798 | 0.5505 | 0.3165 | 0.0771 | 0.2527 | 0.2287 | 0.2074 |
| human_sexuality | 0.5635 | 0.5317 | 0.2302 | 0.0317 | 0.2778 | 0.2143 | 0.3651 |
| agronomy | 0.5621 | 0.5089 | 0.1775 | 0.0651 | 0.2071 | 0.2485 | 0.4201 |
| sports_science | 0.6424 | 0.5697 | 0.2000 | 0.0909 | 0.2848 | 0.2667 | 0.3879 |
| nutrition | 0.5586 | 0.4759 | 0.1862 | 0.0621 | 0.1172 | 0.2414 | 0.2552 |
| elementary_IT | 0.5462 | 0.5378 | 0.2353 | 0.1050 | 0.1975 | 0.2353 | 0.2647 |
For Conditional Accuracy, as shown in Table IX, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead. Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities.
| Category | Conditional Accuracy | ||||||
| S-S | G-4o | L-318 | G-35T | Q-2.5 | Go1-m | DK-r1 | |
| ancient_chinese | 0.2697 | 0.2500 | 0.2174 | 0.2500 | 0.2195 | 0.2812 | 0.2500 |
| chinese_civil_service_exam | 0.2562 | 0.2152 | 0.2222 | 0.1429 | 0.3243 | 0.2368 | 0.2658 |
| chinese_food_culture | 0.3258 | 0.3143 | 0.3529 | 0.2727 | 0.3226 | 0.2143 | 0.3171 |
| chinese_foreign_policy | 0.3444 | 0.3438 | 0.2973 | 0.0000 | 0.2667 | 0.3235 | 0.3448 |
| chinese_teacher_qualification | 0.2921 | 0.2752 | 0.1667 | 0.3077 | 0.3636 | 0.2571 | 0.2650 |
| elementary_chinese | 0.3126 | 0.2813 | 0.2857 | 0.4286 | 0.2958 | 0.2533 | 0.3145 |
| ethnology | 0.3719 | 0.4109 | 0.2400 | 0.2308 | 0.3864 | 0.3333 | 0.5714 |
| high_school_politics | 0.3662 | 0.3225 | 0.2963 | 0.2143 | 0.1667 | 0.2174 | 0.3412 |
| modern_chinese | 0.3592 | 0.3265 | 0.2000 | 0.4000 | 0.1250 | 0.1852 | 0.3489 |
| chinese_literature | 0.2980 | 0.2000 | 0.3030 | 0.1765 | 0.3220 | 0.3077 | 0.2879 |
| chinese_history | 0.3407 | 0.2552 | 0.2778 | 0.2941 | 0.2353 | 0.2418 | 0.2543 |
| chinese_driving_rule | 0.3299 | 0.2764 | 0.1923 | 0.1111 | 0.2424 | 0.3247 | 0.3721 |
| construction_project_management | 0.3470 | 0.2951 | 0.2593 | 0.2222 | 0.2162 | 0.2292 | 0.3253 |
| elementary_commonsense | 0.3017 | 0.2925 | 0.2500 | 0.4667 | 0.1964 | 0.2759 | 0.2654 |
| traditional_chinese_medicine | 0.2971 | 0.2100 | 0.2045 | 0.2308 | 0.3023 | 0.2738 | 0.2912 |
| business_ethics | 0.3981 | 0.3411 | 0.2791 | 0.3333 | 0.2623 | 0.2391 | 0.3802 |
| economics | 0.3597 | 0.3476 | 0.2955 | 0.2857 | 0.1500 | 0.3341 | 0.3254 |
| education | 0.3667 | 0.3421 | 0.3647 | 0.2667 | 0.3111 | 0.1818 | 0.3845 |
| college_education | 0.3125 | 0.2143 | 0.3600 | 0.2353 | 0.2174 | 0.2857 | 0.3478 |
| journalism | 0.2981 | 0.2366 | 0.2667 | 0.2857 | 0.2679 | 0.2281 | 0.3098 |
| marketing | 0.3768 | 0.3051 | 0.2683 | 0.2500 | 0.3750 | 0.3541 | 0.3254 |
| professional_accounting | 0.3748 | 0.3084 | 0.3438 | 0.5714 | 0.2647 | 0.3451 | 0.3746 |
| professional_psychology | 0.3985 | 0.2823 | 0.2157 | 0.3158 | 0.3333 | 0.3698 | 0.3945 |
| public_relations | 0.2427 | 0.1889 | 0.3600 | 0.1333 | 0.2250 | 0.2845 | 0.4752 |
| security_study | 0.2997 | 0.3333 | 0.3182 | 0.3125 | 0.3636 | 0.2258 | 0.3977 |
| high_school_geography | 0.4179 | 0̆.2807 | 0.2188 | 0.1250 | 0.2308 | 0.2368 | 0.3627 |
| management | 0.3089 | 0.2913 | 0.2264 | 0.3333 | 0.2941 | 0.3154 | 0.5369 |
| sociology | 0.3333 | 0.2997 | 0.2889 | 0.1000 | 0.1562 | 0.2062 | 0.3221 |
| marxist_theory | 0.3217 | 0.3143 | 0.2453 | 0.1000 | 0.2712 | 0.3210 | 0.3212 |
| college_law | 0.3287 | 0.2842 | 0.2800 | 0.2222 | 0.2258 | 0.2165 | 0.3109 |
| global_facts | 0.3636 | 0.3599 | 0.3235 | 0.0909 | 0.2353 | 0.3000 | 0.3611 |
| international_law | 0.3830 | 0.2524 | 0.2381 | 0.3636 | 0.2903 | 0.2714 | 0.3799 |
| jurisprudence | 0.3258 | 0.2667 | 0.2519 | 0.2857 | 0.2754 | 0.2786 | 0.3299 |
| world_religions | 0.5975 | 0.2911 | 0.2222 | 0.2667 | 0.2857 | 0.2143 | 0.6475 |
| logical | 0.2931 | 0.2879 | 0.2778 | 0.2500 | 0.2593 | 0.2849 | 0.3648 |
| professional_law | 0.3178 | 0.2223 | 0.1786 | 0.2000 | 0.2609 | 0.2865 | 0.3802 |
| philosophy | 0.3276 | 0.2429 | 0.2308 | 0.3000 | 0.3478 | 0.3169 | 0.3057 |
| world_history | 0.5658 | 0.2907 | 0.2286 | 0.5455 | 0.3148 | 0.3267 | 0.4407 |
| arts | 0.2759 | 0.2500 | 0.3333 | 0.4375 | 0.3684 | 0.2157 | 0.2453 |
| electrical_engineering | 0.3349 | 0.3478 | 0.3114 | 0.1304 | 0.2745 | 0.2456 | 0.2857 |
| engineering_hydrology | 0.3684 | 0.1148 | 0.1304 | 0.5714 | 0.2121 | 0.2500 | 0.4730 |
| astronomy | 0.3903 | 0.2237 | 0.2750 | 0.3750 | 0.3333 | 0.2111 | 0.3699 |
| medical_statistics | 0.2978 | 0.2656 | 0.2222 | 0.5000 | 0.2105 | 0.3636 | 0.3026 |
| computer_science | 0.4013 | 0.2845 | 0.2500 | 0.2778 | 0.3600 | 0.3966 | 0.3923 |
| conceptual_physics | 0.2729 | 0.2473 | 0.2712 | 0.2000 | 0.2174 | 0.2632 | 0.2799 |
| anatomy | 0.2410 | 0.2338 | 0.1333 | 0.2000 | 0.2000 | 0.1667 | 0.2367 |
| machine_learning | 0.2319 | 0.2241 | 0.2083 | 0.1429 | 0.2000 | 0.2498 | 0.2729 |
| high_school_physics | 0.2667 | 0.1857 | 0.1429 | 0.1250 | 0.2593 | 0.2441 | 0.6366 |
| high_school_mathematics | 0.6667 | 0.5000 | 0.3333 | 0.4500 | 0.5268 | 0.5246 | 0.5241 |
| high_school_chemistry | 0.2667 | 0.1974 | 0.3590 | 0.1111 | 0.2857 | 0.2459 | 0.3636 |
| high_school_biology | 0.3171 | 0.2544 | 0.2727 | 0.5000 | 0.2500 | 0.2321 | 0.3295 |
| elementary_mathematics | 0.3146 | 0.2519 | 0.2273 | 0.2778 | 0.4615 | 0.2830 | 0.2941 |
| actuarial_science | 0.3400 | 0.2500 | 0.4444 | 0.1579 | 0.3214 | 0.2800 | 0.4737 |
| genetics | 0.2941 | 0.2555 | 0.3529 | 0.2143 | 0.3793 | 0.3011 | 0.3582 |
| college_mathematics | 0.5882 | 0.2642 | 0.2727 | 0.2000 | 0.4941 | 0.1905 | 0.4896 |
| virology | 0.2083 | 0.1744 | 0.1200 | 0.4286 | 0.2188 | 0.3094 | 0.2157 |
| legal_and_moral_basis | 0.3821 | 0.3277 | 0.2963 | 0.1579 | 0.3284 | 0.2436 | 0.3684 |
| computer_security | 0.3223 | 0.2848 | 0.3171 | 0.1667 | 0.3167 | 0.2750 | 0.3033 |
| food_science | 0.3914 | 0.3824 | 0.1905 | 0.2308 | 0.2703 | 0.2250 | 0.4154 |
| college_medicine | 0.3134 | 0.2500 | 0.2055 | 0.2000 | 0.2128 | 0.2022 | 0.3326 |
| clinical_knowledge | 0.3448 | 0.3088 | 0.2222 | 0.3333 | 0.3426 | 0.2143 | 0.3248 |
| professional_medicine | 0.2431 | 0.2222 | 0.2605 | 0.3448 | 0.2737 | 0.2569 | 0.3433 |
| human_sexuality | 0.3134 | 0.3099 | 0.1726 | 0.7500 | 0.3714 | 0.3174 | 0.2989 |
| agronomy | 0.2791 | 0.2558 | 0.2667 | 0.0909 | 0.3714 | 0.2188 | 0.2626 |
| sports_science | 0.4171 | 0.2666 | 0.3936 | 0.2667 | 0.4465 | 0.2969 | 0.3830 |
| nutrition | 0.2319 | 0.2222 | 0.1481 | 0.2222 | 0.2941 | 0.3833 | 0.4412 |
| elementary_IT | 0.3185 | 0.3906 | 0.2500 | 0.3600 | 0.2128 | 0.2874 | 0.3452 |
For Accuracy, as shown in Table X, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
| Category | Accuracy | ||||||
| S-S | G-4o | L-318 | G-35T | Q-2.5 | Go1-m | DK-r1 | |
| ancient_chinese | 0.1341 | 0.1220 | 0.3050 | 0.0183 | 0.0549 | 0.0549 | 0.1159 |
| chinese_civil_service_exam | 0.1062 | 0.1000 | 0.0375 | 0.0062 | 0.0750 | 0.0025 | 0.1562 |
| chinese_food_culture | 0.1765 | 0.1618 | 0.0441 | 0.0221 | 0.0735 | 0.0294 | 0.1838 |
| chinese_foreign_policy | 0.2056 | 0.2028 | 0.1028 | 0.0000 | 0.0748 | 0.0654 | 0.1869 |
| chinese_teacher_qualification | 0.1676 | 0.1553 | 0.0447 | 0.0223 | 0.0894 | 0.0503 | 0.2458 |
| elementary_chinese | 0.1866 | 0.1695 | 0.0476 | 0.0238 | 0.0833 | 0.0397 | 0.1865 |
| ethnology | 0.2296 | 0.2456 | 0.0444 | 0.0222 | 0.1259 | 0.0519 | 0.3333 |
| high_school_politics | 0.1738 | 0.1834 | 0.0559 | 0.0210 | 0.0350 | 0.0210 | 0.2098 |
| modern_chinese | 0.1897 | 0.1654 | 0.0345 | 0.0345 | 0.0259 | 0.0603 | 0.1466 |
| chinese_literature | 0.1765 | 0.1882 | 0.0490 | 0.0147 | 0.0931 | 0.0686 | 0.1961 |
| chinese_history | 0.1920 | 0.1146 | 0.0464 | 0.0155 | 0.0619 | 0.0372 | 0.1238 |
| chinese_driving_rule | 0.1676 | 0.1492 | 0.0382 | 0.0076 | 0.0611 | 0.0763 | 0.1985 |
| construction_project_management | 0.2230 | 0.1295 | 0.0504 | 0.0144 | 0.0576 | 0.0504 | 0.1727 |
| elementary_commonsense | 0.1966 | 0.1714 | 0.0354 | 0.0354 | 0.0556 | 0.0354 | 0.2424 |
| traditional_chinese_medicine | 0.1568 | 0.1323 | 0.0486 | 0.0162 | 0.0703 | 0.0703 | 0.0973 |
| business_ethics | 0.2057 | 0.1852 | 0.0574 | 0.0191 | 0.0766 | 0.0478 | 0.1914 |
| economics | 0.1195 | 0.1195 | 0.0818 | 0.0126 | 0.0377 | 0.0252 | 0.1824 |
| education | 0.1865 | 0.1795 | 0.0675 | 0.0245 | 0.0859 | 0.0613 | 0.1779 |
| college_education | 0.1867 | 0.2100 | 0.0741 | 0.0374 | 0.0935 | 0.1028 | 0.2056 |
| journalism | 0.1829 | 0.1676 | 0.0465 | 0.0116 | 0.0872 | 0.0640 | 0.2035 |
| marketing | 0.1833 | 0.2000 | 0.0611 | 0.0222 | 0.1000 | 0.1167 | 0.1667 |
| professional_accounting | 0.1886 | 0.1714 | 0.0629 | 0.0457 | 0.0514 | 0.0629 | 0.1523 |
| professional_psychology | 0.1897 | 0.1736 | 0.0474 | 0.0259 | 0.0647 | 0.0690 | 0.1681 |
| public_relations | 0.1437 | 0.0977 | 0.0517 | 0.0115 | 0.0517 | 0.0862 | 0.1389 |
| security_study | 0.1333 | 0.2222 | 0.0519 | 0.0370 | 0.1481 | 0.0593 | 0.1259 |
| high_school_geography | 0.1356 | 0.2333 | 0.0593 | 0.0085 | 0.0763 | 0.0508 | 0.2034 |
| management | 0.1810 | 0.1695 | 0.0571 | 0.0286 | 0.0714 | 0.1163 | 0.2143 |
| sociology | 0.1593 | 0.1726 | 0.0575 | 0.0044 | 0.0442 | 0.0619 | 0.1681 |
| marxist_theory | 0.1957 | 0.1799 | 0.0688 | 0.0053 | 0.0847 | 0.0741 | 0.2222 |
| college_law | 0.1574 | 0.1222 | 0.0648 | 0.0185 | 0.0648 | 0.0833 | 0.1759 |
| global_facts | 0.1822 | 0.1621 | 0.0738 | 0.0067 | 0.0805 | 0.0671 | 0.1409 |
| international_law | 0.1405 | 0.1965 | 0.0541 | 0.0216 | 0.0973 | 0.0595 | 0.0270 |
| jurisprudence | 0.1405 | 0.1998 | 0.0803 | 0.0195 | 0.0925 | 0.0706 | 0.1752 |
| world_religions | 0.1437 | 0.1188 | 0.0250 | 0.0250 | 0.0750 | 0.0312 | 0.2750 |
| logical | 0.1546 | 0.1382 | 0.0407 | 0.0244 | 0.0569 | 0.0813 | 0.1220 |
| professional_law | 0.1611 | 0.1009 | 0.0474 | 0.0095 | 0.0569 | 0.0474 | 0.0758 |
| philosophy | 0.0667 | 0.1810 | 0.0571 | 0.0286 | 0.0762 | 0.0571 | 0.0095 |
| world_history | 0.1543 | 0.1304 | 0.0497 | 0.0373 | 0.1056 | 0.0807 | 0.1577 |
| arts | 0.1880 | 0.1500 | 0.5000 | 0.0438 | 0.0875 | 0.0438 | 0.1750 |
| electrical_engineering | 0.1429 | 0.1190 | 0.0581 | 0.0174 | 0.0814 | 0.0523 | 0.2209 |
| engineering_hydrology | 0.1981 | 0.1799 | 0.0283 | 0.3770 | 0.0660 | 0.0755 | 0.1792 |
| astronomy | 0.1273 | 0.1030 | 0.0667 | 0.0242 | 0.0545 | 0.0970 | 0.1697 |
| medical_statistics | 0.1604 | 0.1415 | 0.0377 | 0.0283 | 0.0377 | 0.0472 | 0.0283 |
| computer_science | 0.1618 | 0.1275 | 0.0049 | 0.0245 | 0.0441 | 0.0441 | 0.1127 |
| conceptual_physics | 0.1497 | 0.1361 | 0.1020 | 0.0068 | 0.0340 | 0.0748 | 0.1224 |
| anatomy | 0.2230 | 0.1216 | 0.0270 | 0.0135 | 0.0338 | 0.0405 | 0.0878 |
| machine_learning | 0.1311 | 0.1000 | 0.0410 | 0.0082 | 0.0328 | 0.0328 | 0.1066 |
| high_school_physics | 0.1818 | 0.1964 | 0.0273 | 0.0091 | 0.0636 | 0.0455 | 0.2091 |
| high_school_mathematics | 0.1182 | 0.1453 | 0.0549 | 0.0549 | 0.1341 | 0.0671 | 0.2378 |
| high_school_chemistry | 0.1136 | 0.0970 | 0.1061 | 0.0076 | 0.0455 | 0.0682 | 0.0303 |
| high_school_biology | 0.1716 | 0.1534 | 0.0710 | 0.0414 | 0.0414 | 0.0651 | 0.0178 |
| elementary_mathematics | 0.1432 | 0.1217 | 0.0435 | 0.0435 | 0.0783 | 0.0478 | 0.1139 |
| actuarial_science | 0.0755 | 0.0472 | 0.0755 | 0.0283 | 0.0849 | 0.0660 | 0.2453 |
| genetics | 0.1307 | 0.1403 | 0.0682 | 0.0170 | 0.0625 | 0.0739 | 0.0625 |
| college_mathematics | 0.1333 | 0.0952 | 0.0571 | 0.0286 | 0.0952 | 0.0286 | 0.0762 |
| virology | 0.1183 | 0.0888 | 0.0178 | 0.0178 | 0.0414 | 0.1183 | 0.0769 |
| legal_and_moral_basis | 0.2182 | 0.2022 | 0.0748 | 0.0140 | 0.1028 | 0.0654 | 0.3411 |
| computer_security | 0.2047 | 0.1579 | 0.0076 | 0.0175 | 0.0877 | 0.0877 | 0.1228 |
| food_science | 0.1366 | 0.1936 | 0.0280 | 0.0210 | 0.0699 | 0.0140 | 0.1329 |
| college_medicine | 0.1538 | 0.1245 | 0.0549 | 0.0073 | 0.0366 | 0.0366 | 0.0623 |
| clinical_knowledge | 0.1772 | 0.1565 | 0.0422 | 0.0338 | 0.0844 | 0.0591 | 0.0549 |
| professional_medicine | 0.1410 | 0.1230 | 0.0824 | 0.0266 | 0.0691 | 0.0718 | 0.0691 |
| human_sexuality | 0.1746 | 0.1667 | 0.0397 | 0.0238 | 0.1032 | 0.0635 | 0.1667 |
| agronomy | 0.1420 | 0.1302 | 0.0473 | 0.0059 | 0.0769 | 0.0533 | 0.2071 |
| sports_science | 0.2242 | 0.2189 | 0.0727 | 0.0242 | 0.1091 | 0.0788 | 0.1333 |
| nutrition | 0.1241 | 0.1103 | 0.0276 | 0.0138 | 0.0345 | 0.0759 | 0.1034 |
| elementary_IT | 0.1092 | 0.2102 | 0.0588 | 0.0378 | 0.0420 | 0.0798 | 0.1387 |
This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient. Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
As shown in Fig. 10, we also compiled statistics on various indicators of Sun-Shine in each sub-branch. Whether it is the best or the second best indicator, Sun-Shine maintains a large advantage in total, whether compared with DeepSeek-R1 or Chat-4o. Especially in the sub-category, “Other”, Sun-Shine obtained the best indicators of all accuracy. The remaining sub-categories also outperform DeepSeek-r1 overall. Only in the indicator ”Conditional Accuracy”, Sun-Shine is slightly lower than the latest DeepSeek-R1.
Also, as shown in Table XI, we also calculated the average of different metrics for different large models on different categories. Sun-Shine is either first or second in all categories and all indicators. Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
| LLM | Response Rate | ||||
| China | Social | Human | STEM | Other | |
| Sun-Shine | 0.5507 | 0.5588 | 0.5377 | 0.5481 | 0.5498 |
| GPT-3.5-turbo | 0.0687 | 0.0741 | 0.0748 | 0.0806 | 0.0801 |
| GPT-4o | 0.4932 | 0.5133 | 0.5107 | 0.4915 | 0.4957 |
| GPT-o1-mini | 0.2050 | 0.2340 | 0.2333 | 0.2231 | 0.2241 |
| Llama-3.18B | 0.1907 | 0.1998 | 0.2056 | 0.2082 | 0.2104 |
| Qwen-2.5 | 0.2563 | 0.2703 | 0.2736 | 0.2441 | 0.2419 |
| DeepSeek-R1 | 0.4206 | 0.4069 | 0.3884 | 0.3580 | 0.3482 |
| LLM | Conditional Accuracy | ||||
| China | Social | Human | STEM | Other | |
| Sun-Shine | 0.3208 | 0.3367 | 0.3477 | 0.3457 | 0.3489 |
| GPT-3.5-turbo | 0.2489 | 0.2605 | 0.2656 | 0.2710 | 0.2731 |
| GPT-4o | 0.2845 | 0.2871 | 0.2846 | 0.2746 | 0.2775 |
| GPT-o1-mini | 0.2636 | 0.2700 | 0.2716 | 0.2742 | 0.2728 |
| Llama-3.18B | 0.2510 | 0.2704 | 0.2662 | 0.2620 | 0.2597 |
| Qwen-2.5 | 0.2656 | 0.2655 | 0.2710 | 0.2866 | 0.2900 |
| DeepSeek-R1 | 0.3209 | 0.3482 | 0.3524 | 0.3584 | 0.3573 |
| LLM | Accuracy | ||||
| China | Social | Human | STEM | Other | |
| Sun-Shine | 0.1888 | 0.1742 | 0.1683 | 0.1614 | 0.1618 |
| GPT-3.5-turbo | 0.0182 | 0.0201 | 0.0205 | 0.0277 | 0.0265 |
| GPT-4o | 0.1594 | 0.1676 | 0.1634 | 0.1515 | 0.1535 |
| GPT-o1-mini | 0.0475 | 0.0584 | 0.0598 | 0.0603 | 0.0606 |
| Llama-3.18B | 0.0656 | 0.0628 | 0.0723 | 0.0662 | 0.0663 |
| Qwen-2.5 | 0.0691 | 0.0723 | 0.0744 | 0.0702 | 0.0709 |
| DeepSeek-R1 | 0.1863 | 0.1819 | 0.1705 | 0.1562 | 0.1534 |
Ti-Safetybench: as shown in Fig. 11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
As shown in Table XII, it presents a comprehensive comparison of leading LLMs on the Ti-Safetybench, evaluating their performance across key safety-related categories. Notably, Sun-Shine outperforms all baseline models under both DA and CCA tasks. In the DA setting, Sun-Shine achieves the highest exact match (EM) score (0.292) and leads in IA, OFF, and UB, while maintaining a significantly higher filter rate (59.20%) than GPT-4o (38.59%) and others. In the CCA setting, Sun-Shine demonstrates even stronger results, ranking first across nearly all categories, including EM (0.483), IA (0.455), mental health (MH, 0.507), and UB (0.466). These results highlight Sun-Shine’s superior ability to generate safe, accurate, and instruction-aligned responses, setting a new benchmark in safety-aware language modeling.
| LLM | Category | ||||||||
| Filter | Invaild | EM | IA | MH | OFF | PH | PP | UB | |
| Sun-Shine | DA | 59.20% | 0.292 | 0.379 | 0.204 | 0.255 | 0.258 | 0.307 | 0.372 |
| GPT-4o | 38.59% | 0.144 | 0.357 | 0.241 | 0.199 | 0.246 | 0.296 | 0.358 | |
| Llama-3.18B | 74.75% | 0.105 | 0.088 | 0.101 | 0.097 | 0.094 | 0.099 | 0.109 | |
| GPT-3.5-turbo | 73.42% | 0.085 | 0.073 | 0.087 | 0.086 | 0.097 | 0.106 | 0.123 | |
| Qwen-2.5 | 49.50% | 0.199 | 0.118 | 0.164 | 0.198 | 0.150 | 0.159 | 0.371 | |
| GPT-o1-mini | 71.20% | 0.109 | 0.070 | 0.095 | 0.111 | 0.097 | 0.095 | 0.170 | |
| DeepSeek-R1 | 51.14% | 0.287 | 0.373 | 0.232 | 0.250 | 0.297 | 0.309 | 0.371 | |
| Sun-Shine | CCA | 16.60% | 0.483 | 0.455 | 0.507 | 0.448 | 0.509 | 0.336 | 0.466 |
| GPT-4o | 34.55% | 0.478 | 0.429 | 0.429 | 0.287 | 0.255 | 0.387 | 0.401 | |
| Llama-3.18B | 69.72% | 0.130 | 0.101 | 0.123 | 0.124 | 0.119 | 0.115 | 0.126 | |
| GPT-3.5-turbo | 67.75% | 0.108 | 0.082 | 0.103 | 0.125 | 0.114 | 0.118 | 0.160 | |
| Qwen-2.5 | 16.61% | 0.448 | 0.413 | 0.427 | 0.289 | 0.259 | 0.385 | 0.403 | |
| GPT-o1-mini | 61.69% | 0.254 | 0.099 | 0.134 | 0.163 | 0.135 | 0.117 | 0.231 | |
| DeepSeek-R1 | 28.14% | 0.558 | 0.457 | 0.500 | 0.429 | 0.511 | 0.339 | 0.457 | |
Overall: Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
5.2 Sun-Shine for Tibetan Culture
As shown in Fig. 12, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
As shown in Fig. 13(a), Fig. 13(b), and Fig. 13(c), we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English.
Sun-Shine generates outputs that are notably neutral, emotionally grounded, and culturally restrained across all three languages. The Tibetan response is respectful and sincere, avoiding ornate metaphors or religious references. In Chinese, the model articulates gratitude with clarity and warmth, emphasizing familial support without exaggeration. The English version maintains a reflective tone, expressing appreciation without invoking religious or hyperbolic imagery. This demonstrates Sun-Shine’s alignment with real-world communication norms and its ability to generalize across linguistic and cultural contexts.
In contrast, GPT-4o exhibits a distinctly elevated and ceremonial tone. Its Tibetan and Chinese outputs are imbued with grandiloquent phrases and spiritual imagery. For example, the model refers to the father as a “beacon of happiness” and ”born on the edge of happiness,” which, while poetic, may not align with typical expressions of filial appreciation. The English version similarly includes metaphor-heavy language, such as “the victory banner is my prayer,” which resembles religious scripture more than natural expression. This suggests a stylistic bias toward exalted, sometimes deific, portrayals.
DeepSeek-R1, on the other hand, appears to have an even stronger inclination toward religious and cultural iconography. It references “Mañjuśrī’s Father” and constructs its Tibetan output with vocabulary and structure closely mirroring classical Buddhist literature. The Chinese response is elegant yet heavily stylized, and the English output reads like a translated verse from scripture rather than a naturally composed poem. While aesthetically refined, this stylistic preference may reflect overfitting to religious texts in training data, limiting the model’s adaptability to general-purpose use.
In summary, Sun-Shine’s responses are more balanced and less prone to embellishment, demonstrating cultural sensitivity and instruction alignment. GPT-4o and DeepSeek-R1, while fluent and creative, tend to overgenerate in stylistically loaded contexts, risking a mismatch with the intended pragmatic tone. This comparison underscores Sun-Shine’s strength in producing appropriate, grounded, and respectful language across multiple languages and cultural domains.
5.3 Downstream Task
Tibetan Language Proficiency Test: we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
| LLM | A | B | C | D |
| Sun-Shine | 63 | 61 | 44 | 42 |
| GPT-4o | 45 | 47 | 53 | 51 |
| DeepSeek-R1 | 59 | 56 | 67 | 66 |
| Qwen-2.5 | 41 | 43 | 53 | 54 |
As shown in Table XIII, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our benchmark, TIB-STC, which includes a richer collection of ancient Tibetan texts.
Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU [40] and ROUGE [41] are used to evaluate this task. As shown in Table XIV, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 12 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
| LLM | B-4 | R-1 | R-2 | R-l |
| Sun-Shine | 42.16 | 65.78 | 48.91 | 51.58 |
| GPT-4o | 38.85 | 61.12 | 44.33 | 47.02 |
| DeepSeek-R1 | 41.85 | 65.49 | 47.73 | 51.04 |
| Qwen-2.5 | 35.20 | 58.59 | 41.25 | 44.73 |
6 Conclusion
In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC benchmark, marking significant progress in Tibetan language processing. Sun-Shine, trained on the comprehensive TIB-STC benchmark, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios. TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data. Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan. Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our benchmark, TIB-STC, in providing a high-quality ancient Tibetan corpus. This work advances low-resource language models and underscores the importance of culturally specific benchmarks.
Acknowledgment
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.

### v4

@@
-Sun-Shine: A Foundation Large Language Model for Tibetan Culture and Heritage
+TIB-STC: A Large-Scale Structured Tibetan Benchmark for Low-Resource Language Modeling
 Abstract
-Tibetan, a minority language in China, features a highly intricate grammatical structure, characterized by four verb tenses and a tense system with frequent irregularities, contributing to its extensive inflectional diversity.
-Recently, advances in Large Language Models (LLMs) have transformed the paradigm in many domains.
-Despite the success in other fields, current LLMs often fall short in catering to the needs of domain experts like Tibetans, and the potential of LLMs for Tibetan culture is under-explored.
-The intrinsic reasons are the immense and intricate nature of Tibetan culture as well as the necessity for higher granularity and richness in knowledge.
-Simultaneously, the complexity and uniqueness of its grammatical structure, coupled with its status as a minority ethnic language, contribute to data scarcity, which remains a fundamental challenge.
-To alleviate these issues, we introduce Llama-Sunshine (Sun-Shine), the first large language model for Tibetan culture, which is expert in various Tibetan language processing tasks.
-Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features.
-We also propose TIB-STC, a comprehensive benchmark comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale benchmark for Tibetan culture.
-Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis.
-Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities.
-Index Terms:
-Large Language Model, Foundation Model, Tibetan Language Benchmark, Tibetan Language Processing1 Introduction
-Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT [1], Gemini [12], Qwen-2.5 [34, 33], Llama [2, 20] and DeepSeek [37, 38].
-Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges [4].
-Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA [2] and DeepSeek [37, 38], which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
-Despite their many strengths, LLMs are not tailored to specific fields.
-Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain [3, 5].
-Also, there are some biases against LLMs [6, 7, 8].
-LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage [4].
-For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments [9].
-Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability.
-Overall, data scarity is the key issue for these areas.
-Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem [11].
-In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language [10].
-Assembling a benchmark about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort.
-Only then can we develop a more specialized and unbiased LLM model on this basis.
-To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus benchmark designed to address the issue of data scarcity in Tibetan language research.
-This benchmark spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan benchmark in the world.
-On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture.
-the first LLM specifically tailored for Tibetan language processing.
-During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
-Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 [38] totally, showing a higher level of knowledge expertise in Tibetan Culture.
-Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture.
-Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and benchmark for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
-In summary, our contributions can be summarized as follows:
-We present Sun-Shine, the first LLM for Tibetan culture, leveraging Chinese-to-Tibetan translation and manual annotation to tackle linguistic complexities and enhance training accuracy, paving the way for NLP in low-resource languages.
-We introduce TIB-STC, the first large-scale Tibetan benchmark, addressing data scarcity in Tibetan language research.
-Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
-Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our benchmark, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
-2 Related Work
-2.1 Tibetan Language Processing
-LLMs for rare or low-resource languages are relatively uncommon [13].
-The absence of corpus data is the primary and most critical issue [9, 11] like Tibetan.
-However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently [16].
-As a result, digital Tibetan language resources are exceptionally scarce.
-This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing [11, 10, 26, 19].
-For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Fig. 1.
-As shown in Fig.
-2, the native Tibetan syllable structure, includes seven kinds of letters.
-There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel.
-Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data.
-To solve this issue, we propose the first large-scale Tibetan benchmark, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
-2.2 Large Language Model
-LLMs, including GPT [4], Gemini [12], and ChatGLM [6, 7], are highly parameterized models with billions of parameters, pre-trained on massive natural language benchmarks [14].
-Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities.
-By applying task-specific instructions and fine-tuning techniques [15, 17, 18], LLMs can be optimized for specialized fields, enhancing their performance and domain expertise [5].
-Taking into account the unique characteristics of Tibetan culture and in order to make life more convenient for the Tibetan people, we introduce the first LLM for Tibetan, Sun-Shine which is based on TIB-STC and cover a wide range of subjects, such as medicine, law, sports, daily life and basic sciences, providing adequate support can significantly improve the daily lives of people in Tibetan areas.
-3 TIB-STC and Sun-Shine
-3.1 Benchmark: TIB-STC
-TIB-STC, is the first comprehensive and detailed Tibetan corpus benchmark for LLM, consisting of five categories of data as shown in the benchmark branch of Table I.
+Advancement of large language models (LLMs) has brought transformative capabilities to NLP, but such progress remains unevenly distributed, especially for low-resource and culturally rich languages like Tibetan.
+In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain benchmark specifically designed to support the development and evaluation of LLMs for the Tibetan language.
+Spanning over 11 billion tokens across literature, religion, medicine, law, and daily communication, TIB-STC preserves traditional grammar and stylistic richness.
+To validate its utility, we train a reference model, Sun-Shine, on TIB-STC through a three-stage pipeline involving pretraining, supervised fine-tuning, and preference optimization.
+Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the benchmark’s effectiveness in enabling robust instruction-following and culturally aligned generation.
+We release TIB-STC to advance research in low-resource language modeling and promote inclusivity in multilingual NLP.
+All data are available: https://github.com/Vicentvankor/sun-shine.
+Introduction
+Large language models (LLMs) have demonstrated impressive performance across a broad range of natural language processing (NLP) tasks, exhibiting strong generalization and emergent capabilities (Minaee et al.
+2024).
+However, these advancements have primarily benefited high-resource languages such as English and Chinese (Achiam et al.
+2023; Yang and et al.
+2024; Touvron et al.
+2023), while low-resource and linguistically complex languages—such as Tibetan—remain significantly underserved (Liang et al.
+2024).
+Tibetan presents unique challenges for NLP, including complex morphological structures, vertically composed syllables, and deep integration with cultural semantics (UCLA International Institute 2014; Khysru et al.
+2021).
+These features, combined with a severe lack of large-scale, domain-diverse, and annotated corpora, have hindered the development of effective LLMs for Tibetan (Gao et al.
+2025).
+Moreover, mainstream LLMs are typically trained on corpora heavily biased toward dominant languages and cultures (Üstün et al.
+2024), limiting their generalization to minority languages.
+As a result, they often produce outputs that are inaccurate, unsafe, or culturally inappropriate when applied to underrepresented languages (Chao et al.
+2024).
+This underscores the urgent need for high-quality, task-aligned benchmarks that support robust and culturally informed modeling for low-resource languages (Brown et al.
+2020; Blodgett et al.
+2020).
+To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus benchmark.
+Covering over 11 billion tokens from domains such as literature, law, religion, medicine, and daily communication, TIB-STC is designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
+To validate the benchmark’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline.
+While the primary contribution of this work lies in the construction of the TIB-STC dataset, Sun-Shine serves as a proof of concept for building culturally aligned and linguistically robust LLMs for Tibetan.
+Our main contributions are summarized as follows:
+We introduce TIB-STC, the first large-scale and structured Tibetan corpus benchmark, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
+We have constructed four task-specific sub-benchmarks (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
+We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the benchmark’s effectiveness in instruction-following and safety-aware generation.
+We evaluate our approach on the TLUE benchmark, a Tibetan adaptation of multilingual evaluation datasets, manually corrected for linguistic and cultural fidelity, highlighting the feasibility of structured benchmarking for minority languages.
+Related Work
+Tibetan Language Processing
+Tibetan remains one of the least-resourced languages in NLP due to its complex grammatical structure, two-dimensional syllabic composition, and limited digitized resources (Liu et al.
+2024).
+The scarcity of Tibetan corpora has severely hindered the development of both traditional NLP tools and modern deep learning models (Liu et al.
+2025a, b).
+Existing work in Tibetan NLP has primarily focused on morphological segmentation, part-of-speech tagging, and rule-based translation, often relying on small, domain-specific datasets or handcrafted features (Liang et al.
+2024).
+However, these approaches lack scalability and generalizability, particularly in the context of large-scale pretraining.
+Moreover, the diverse functional domains of Tibetan, spanning religious scripture, legal and medical texts, and everyday communication, necessitate a corpus that captures both linguistic richness and cultural depth (Khysru et al.
+2021).
+To date, no large-scale, structured benchmark exists to systematically support foundation model training or evaluation for Tibetan.
+Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan benchmark designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
+Large Language Models
+LLMs such as GPT-4 (Achiam et al.
+2023), LLaMA (Touvron et al.
+2023; Dubey et al.
+2024), and DeepSeek (Bi and et al.
+2024; Guo and et al.
+2025) have demonstrated impressive zero-shot and few-shot performance on a variety of tasks in high-resource languages.
+However, their effectiveness diminishes considerably in low-resource or culturally divergent settings.
+Recent efforts have explored domain- or region-specific LLMs by integrating language-specific corpora and instruction-tuning datasets (Taori et al.
+2023b; GLM et al.
+2024; Wang et al.
+2023), but most of these focus on widely spoken or contemporary languages.
+Research on LLMs for minority languages with non-Latin scripts and typologically distinct grammars, such as Tibetan, remains extremely limited (Üstün et al.
+2024; Minaee et al.
+2024).
+In addition, safety alignment and bias mitigation for underrepresented languages are underexplored, despite the increased risk of harmful, inaccurate, or culturally insensitive outputs.
+To address these limitations, we present Sun-Shine, a Tibetan LLM trained on TIB-STC, to empirically validate the effectiveness of culturally grounded data in enabling robust and responsible modeling for low-resource language contexts.
+TIB-STC
+TIB-STC, the first comprehensive and detailed Tibetan corpus benchmark for LLM, consists of five categories of data as shown in the benchmark branch of Table 1.
+All data selection and cleaning are conducted by experts in the Tibetan language.
+Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
+For the TIB-STC, Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement.
+For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
 | Benchmark | Instance | Tokens | Storage Size |
@@
 | hh-rlhf-Ti | 47,411 | 98M | 154MB |
-Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement.
-For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability.
-Pre-Training involves the largest benchmark in terms of instances, tokens, and storage, as it focuses on general language model training.
-In the stage of Supervised Fine-Tuning (SFT), benchmarks are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety).
-Direct Preference Optimization (DPO) uses the smallest benchmarks but focuses on high-quality preference data for alignment and optimization.
-For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language.
-Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
-3.1.1 Corpus
+Corpus
 Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%).
@@
 There are 122,756,184 such instances.
-Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation.
-Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base.
-As shown in Fig.
-3, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language.
+As shown in Figure 1, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language.
 Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
-3.1.2 Alpaca-Ti
+Alpaca-Ti
 Alpaca-Ti is used for instruction fine-tuning.
-We draw inspiration from Alpaca [31, 32] project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world.
-It covers the following topics as shown in Table II.
+We draw inspiration from the Alpaca (Taori et al.
+2023b, a) project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world.
+It covers the following topics as shown in Table 2.
 | Domain | Category |
@@
 It has a total of 42,676 instances.
-One data sample is shown in Fig.
-4.
-3.1.3 Safety-Prompts-Ti
-Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table III, and the data size is 38,156.
-LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts [22, 23, 24].
-| Instruction Attacks | 15k |
-| Instruction Attacks consists of Goal Hijacking | 2.5K |
-| Prompt Leaking | 2.5K |
-| RolePlay Instruction | 2.5K |
-| Unsafe Instruction Topic | 2.5K |
-| Inquiry with Unsafe Opinion | 2.5K |
-| Reverse Exposure | 2.5K |
-| Typical Safety Scenarios | 23K |
-| Typical Safety Scenarios consists of Insult | 3.6K |
-| Unfairness and Discrimination | 3.8K |
-| Crimes and Illegal Activities | 3.6K |
-| Physical Harm | 1.7K |
-| Mental Health | 4K |
-| Privacy and Property | 1.3K |
-| Ethics and Morality | 1.3K |
-| Others | 4.7K |
-Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Fig. 5.
-3.1.4 CValues-Ti
-CValues-Ti is the benchmark, used to improve the security of Tibetan language models.
-Considering work [25, 27, 28], we propose this branch of a small benchmark because LLM can pose risks or has negative social impacts.
-For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response).
+One data sample is shown in Figure 2.
+Safety-Prompts-Ti
+Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios and the data size is 38,156.
+We categorize a total of 15,000 Instruction Attacks into six subtypes, including Goal Hijacking, Prompt Leaking, RolePlay Instructions, Unsafe Instruction Topics, Inquiries with Unsafe Opinions, and Reverse Exposure—each contributing 2,500 instances.
+In addition, we compile 23,000 samples under Typical Safety Scenarios, covering eight safety-critical topics: Insult (3.6K), Unfairness and Discrimination (3.8K), Crimes and Illegal Activities (3.6K), Physical Harm (1.7K), Mental Health (4K), Privacy and Property (1.3K), Ethics and Morality (1.3K), and Others (4.7K).
+LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts (Chao et al.
+2024; Zheng et al.
+2024; Sun et al.
+2023).
+workload it entails, we opted to create a benchmark to simulate the process in advance.
+One data sample is shown in Figure 3.
+CValues-Ti
+CValues-Ti is the benchmark to improve the security of Tibetan language models.
+Considering work (Moore, Deshpande, and Yang 2024; Rozen et al.
+2024; Xu et al.
+2023), we propose this branch of a small benchmark because LLM can pose risks or have negative social impacts.
+For CValues-Ti, it has three categories (safe & responsibility = rejection & positive advice; safe = rejection; unsafe = risk response).
 Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe).
-Unlike the work [28], We streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments.
-One data sample is shown in Fig.
-6.
-3.1.5 hh-rlhf-Ti
+Unlike the work (Xu et al.
+2023), we streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments.
+One data sample is shown in Figure 4.
+hh-rlhf-Ti
 hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention.
-It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF [29, 30], but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives.
+It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF (Bai et al.
+2022; Kirk et al.
+2024), but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives.
 This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan.
-It has a total of 47,411 instances, and one data sample is shown in Fig.
-8.
-3.2 Sun-Shine
-Sun-Shine is built on the LLaMA-3 [2, 20] architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community.
-For more accessible training, we adopt the LLaMA 3.1 (8B) model [20].
-Additionally, we integrate NLP model translation with human annotation [21] as supplementary data to enhance Sun-Shine’s accuracy.
-As illustrated in Fig.
-7, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO).
-For training details, Table IV has shown parameters’ setting.
-Notably, we adopted LoRA-based parameter-efficient fine-tuning [15] for both SFT and DPO, which significantly reduced memory usage without sacrificing performance.
+It has a total of 47,411 instances, and one data sample is shown in Figure 5.
+Sun-Shine
+Sun-Shine is built on the LLaMA-3 (Touvron et al.
+2023; Dubey et al.
+2024) architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community.
+For more accessible training, we adopt the LLaMA 3.1 (8B) model (Dubey et al.
+2024).
+Additionally, we integrate NLP model translation with human annotation (Üstün et al.
+2024) as supplementary data to enhance Sun-Shine’s accuracy.
+As illustrated in Figure 6, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT) (Yu et al.
+2021), and Direct Preference Optimization (DPO) (Rafailov et al.
+2024).
+For training details, Table 3 shows the parameters’ settings.
+Notably, we adopted LoRA-based parameter-efficient fine-tuning (Hu et al.
+2021) for both SFT and DPO, which significantly reduced memory usage without sacrificing performance.
 Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages.
@@
 | pref_loss | – | – | Sigmoid |
-3.2.1 Pre-Training
-During the Pre-training stage, the Sun-Shine is pre-trained on Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction.
+Pre-Training
+During the Pre-training stage, the Sun-Shine is pre-trained on the Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction.
 And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text.
 For details,
-3.2.2 Supervised Fine-Tuning
-Following pretraining, the model undergoes supervised fine-tuning [18] with Alpaca-Ti and Safety-Prompts-Ti.
-During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
-3.2.3 Direct Preference Optimization
-In the stage of DPO [17, 22, 23], the model’s alignment and output quality are further optimized using preference data.
+Supervised Fine-Tuning
+Following pretraining, the model undergoes supervised fine-tuning (Yu et al.
+2021) with Alpaca-Ti and Safety-Prompts-Ti.
+During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, to align the model’s behavior with particular tasks or domains while ensuring safety and professionalism.
+Direct Preference Optimization
+In the stage of DPO (Rafailov et al.
+2024; Chao et al.
+2024; Zheng et al.
+2024), the model’s alignment and output quality are further optimized using preference data.
 The input data comes from CValues-Ti and hh-rlhf-Ti.
@@
 The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
-4 Evaluation: TLUE Benchmark
-Unfortunately, existing benchmarks like CMMLU [35] and Safetybench [36, 24] are available only in Chinese or English, with no purely Tibetan versions.
-To address this issue, we used TLUE [42], the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench.
+Experimental Result
+Evaluation: TLUE Benchmark
+We used TLUE (Gao et al.
+2025), the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench.
 Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
-4.1 Ti-MMLU
-CMMLU [35] is a well-established Chinese benchmark that covers multiple disciplines, including natural sciences, social sciences, engineering, and the humanities.
-To adapt this benchmark for Tibetan, TLUE [42] leveraged an extended version, Ti-MMLU, where the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance.
+Ti-MMLU
+Ti-MMLU is based on CMMLU (Li et al.
+2024), and the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance.
 It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions.
 These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions.
-The benchmark includes 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains.
-Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere.
-The evaluation framework follows the original CMMLU metrics, including Response Rate, Accuracy, and Conditional Accuracy.
+The benchmark includes 17 STEM tasks, 13 humanities (Human) tasks, 22 social science (Social) tasks, and 15 tasks from other domains.
+Notably, 16 tasks are China-specific (China), meaning they either lack equivalents in other regions or may have contextually different answers elsewhere.
+The evaluation framework follows the original CMMLU metrics, including Response Rate (RR), Accuracy (ACC), and Conditional Accuracy (CA).
 Response rate measures the proportion of instances where LLMs provide a valid response.
 Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
-4.2 Ti-Safetybench
-SafetyBench [36, 24] is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, and Privacy and Property.
+Ti-Safetybench
+SafetyBench (Zhang and et al.
+2024; Sun et al.
+2023) is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness (OFF), Unfairness and Bias (UB), Physical Health (PH), Mental Health (MH), Illegal Activities (IA), Ethics and Morality (EM), and Privacy and Property (PP).
 Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples.
-To evaluate safety-related capabilities for Tibetan, TLUE [42] employed Ti-Safetybench, a Tibetan adaptation of SafetyBench.
+To evaluate safety-related capabilities for Tibetan, TLUE (Gao et al.
+2025) employed Ti-Safetybench, a Tibetan adaptation of SafetyBench.
 The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards.
-The evaluation methodology incorporates two scoring methods: Direct Answer Calculation and Concern All Answer Calculation.
+The evaluation methodology incorporates two scoring methods: Direct Answer Calculation (DA) and Concern All Answer Calculation (CAA).
 DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations.
 Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
-Note: as shown in Table V, in order to make the paper more aesthetically pleasing and easier for researchers to read, some specialized terms are abbreviated.
-| Category | Full Name | Abbreviation |
-| STEM | STEM | |
-| Humanities | Human | |
-| Ti-MMLU | Social Sciences | Social |
-| Other | Other | |
-| China Specific | China | |
-| Offensiveness | OFF | |
-| Unfairness and Bias | UB | |
-| Physical Health | PH | |
-| Ti-Safetybench | Mental Health | MH |
-| Illegal Activities | IA | |
-| Ethics and Morality | EM | |
-| Privacy and Property | PP | |
-| Response Rate | RR | |
-| Accuracy | ACC | |
-| Evaluation Metrics | Conditional Accuracy | CA |
-| Direct Answer Calculation | DA | |
-| Concern All Answer Calculation | CAA |
-5 Experimental Result
-We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) [4], Llama-3.18B [20, 2], Qwen-2.5 (72B) [34] and DeepSeek-R1 [37, 38].
-Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis.
-As shown in Table VI, regarding open-source model execution, the Qwen-2.5 and DeepSeek-r1 were evaluated via their respective official APIs provided by the model developers.
-The LlaMA-3.18B was accessed and run using the LlaMA-API platform.
-These details will also be clearly documented in the final version to enhance transparency and reproducibility.
+Baseline Model
+We compare Sun-Shine with several widely used LLMs, including GPT-4o, GPT-4o-mini, and GPT-3.5-turbo (Achiam et al.
+2023), LLaMA-3 18B (Dubey et al.
+2024; Touvron et al.
+2023), Qwen-2.5 72B (Yang and et al.
+2024), and DeepSeek-R1 (Bi and et al.
+2024; Guo and et al.
+2025).
 | LLM | Temperature | Top_p | Stream |
@@
 | GPT-o1-mini | 1.0 | 1.0 | True |
-| DeepSeek-R1 | 1.0 | None | Rtue |
+| DeepSeek-R1 | 1.0 | None | True |
 | LlaMA-3.18B | 0.6 | 0.9 | False |
 | Qwen-2.5 | 0.7 | 0.8 | False |
-Also, as shown in Table VII, LLM and some indicators of downstream tasks, such as BLUE [40] and ROUGH [41], used to evaluate translation tasks, also uses abbreviations in some tables for easier reading.
-For the experimental metrics, the best metric is bolded and the suboptimal metric is underlined.
-| Category | Full Name | Abbreviation |
-| LLM | GPT-3.5-turbo | G-35T |
-| GPT-4o | G-4o | |
-| GPT-o1-mini | Go1-m | |
-| DeepSeek-R1 | DK-r1 | |
-| LlaMA-3.18B | L-318B | |
-| Qwen-2.5 | Q-2.5 | |
-| Sun-Shine | S-S | |
-| Evaluation Metrics | BLEU-4 | B-4 |
-| ROUGE-1 | R-1 | |
-| ROUGE-2 | R-2 | |
-| ROUGE-l | R-l |
-5.1 Insights from Performance Results
-Ti-MMLU: as shown in Fig. 9, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs.
-For Response Rate, as shown in Table VIII, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1.
-Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training benchmark.
+Detailed experimental settings and evaluation metrics for each model are provided in Table 4.
+Specifically, Qwen-2.5 and DeepSeek-R1 were tested via their official APIs, while LLaMA-3 18B was accessed through the LLaMA-API platform.
+All experimental configurations will be fully documented in the final version to ensure transparency and reproducibility.
+Insights from Performance Results
+Ti-MMLU
+As shown in Figure 7, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs.
+Among them, Sun-Shine is slightly ahead of these.
+For the Response Rate, Sun-Shine and GPT-4o achieve the highest Response Rate, with Sun-Shine leading across all sub-branches,shown in Figure 8, demonstrating superior responsiveness to Tibetan prompts enabled by its domain-specific training.
+For Conditional Accuracy, Sun-Shine and DeepSeek-R1 perform best, with the latter holding a slight advantage.
+Both models maintain consistently high scores, reflecting strong general knowledge capabilities.
+In terms of overall Accuracy, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
+This underscores that Sun-Shine’s strong Tibetan understanding stems not only from model design but also from effective use of the TIB-STC dataset and targeted training, allowing an 8B model to surpass the 671B MoE DeepSeek-R1 in key tasks.
+Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance.
+Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine.
+For more details, please refer to supplementary materials.
+Ti-Safetybench
+as shown in Figure 9, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods.
+Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA.
+GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
+Overall, Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1.
+While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance.
+This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
+Sun-Shine for Tibetan Culture
+As shown in Figure 10, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese.
+This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records.
+The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
+As shown in Figure 11, we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English.
+Sun-Shine maintains a neutral tone when addressing common questions, avoiding exaggerated praise or references to Tibetan classics.
+In contrast, GPT-4o adopts a more amplified and majestic style, reminiscent of divine scripture, while DeepSeek-R1’s responses frequently allude to the bodhisattva Mañjuśrī, appearing slightly ”overfitted” with more refined yet noticeably embellished language than GPT-4o.
+Overall, Sun-Shine surpasses DeepSeek-R1 and GPT-4o in most performance indicators, with its classical Tibetan style attributed to the curated TIB-STC and training strategy, where human annotation and correction played a key role.
+Downstream Task
+Translation
+we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan).
+BELU (Papineni et al.
+2002) and ROUGE (Lin 2004) are used to evaluate this task.
+As shown in Table 5, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower.
+Table 10 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
+| LLM | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-l |
+| Sun-Shine | 42.16 | 65.78 | 48.91 | 51.58 |
+| GPT-4o | 38.85 | 61.12 | 44.33 | 47.02 |
+| DeepSeek-R1 | 41.85 | 65.49 | 47.73 | 51.04 |
+| Qwen-2.5 | 35.20 | 58.59 | 41.25 | 44.73 |
+Tibetan Language Proficiency Test
+we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
+As shown in Table 6, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics.
+This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our dataset, TIB-STC, which includes a richer collection of ancient Tibetan texts.
+| LLM | A | B | C | D |
+| Sun-Shine | 63 | 61 | 44 | 42 |
+| GPT-4o | 45 | 47 | 53 | 51 |
+| DeepSeek-R1 | 59 | 56 | 67 | 66 |
+| Qwen-2.5 | 41 | 43 | 53 | 54 |
+Conclusion
+In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC benchmark and the Sun-Shine large language model.
+Our primary focus is the development of TIB-STC, the first large-scale, high-quality Tibetan corpus designed to address the severe lack of resources in low-resource language research.
+Covering diverse domains such as literature, religion, news, and dialogue, TIB-STC offers structured, culturally rich data that enables the effective training and evaluation of Tibetan language models.
+Leveraging this benchmark, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression.
+Extensive evaluation shows that models trained on TIB-STC outperform existing LLMs across metrics such as accuracy, response rate, and stylistic alignment.
+These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded benchmarks in advancing low-resource language technologies.
+Future Work
+As part of our ongoing efforts, we have launched Sun-Shine 2.0, an upgraded version of our Tibetan foundation model with enhanced reasoning capabilities.
+In parallel, we are curating an extended Tibetan corpus to further improve linguistic and cultural coverage.
+Acknowledgement
+This work was supported in part by the National Natural Science Foundation of China under Grants 62276055 and 62406062, in part by the Sichuan Science and Technology Program under Grant 2023YFG0288, in part by the Natural Science Foundation of Sichuan Province under Grant 2024NSFSC1476, in part by the National Science and Technology Major Project under Grant 2022ZD0116100, and in part by the Sichuan Provincial Major Science and Technology Project under Grant 2024ZDZX0012.
+Appendix A Supplementary Material
+Note: S-S = Sun-Shine; G-4o = GPT-4o; L-318B = Llama-3.18B; G-35T = GPT-3.5-Turbo; Q-2.5 = Qwen-2.5; Go1-m = GPT-o1-mini; DK-r1 = DeepSeek-r1.
+Details of Response Rate
 | Category | Response Rate | ||||||
@@
 | elementary_IT | 0.5462 | 0.5378 | 0.2353 | 0.1050 | 0.1975 | 0.2353 | 0.2647 |
-For Conditional Accuracy, as shown in Table IX, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead.
-Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities.
+Details of Conditional Accuracy
 | Category | Conditional Accuracy | ||||||
@@
 | elementary_IT | 0.3185 | 0.3906 | 0.2500 | 0.3600 | 0.2128 | 0.2874 | 0.3452 |
-For Accuracy, as shown in Table X, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
+Details of Accuracy
 | Category | Accuracy | ||||||
@@
 | elementary_IT | 0.1092 | 0.2102 | 0.0588 | 0.0378 | 0.0420 | 0.0798 | 0.1387 |
-This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient.
-Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
-As shown in Fig.
-10, we also compiled statistics on various indicators of Sun-Shine in each sub-branch.
-Whether it is the best or the second best indicator, Sun-Shine maintains a large advantage in total, whether compared with DeepSeek-R1 or Chat-4o.
-Especially in the sub-category, “Other”, Sun-Shine obtained the best indicators of all accuracy.
-The remaining sub-categories also outperform DeepSeek-r1 overall.
-Only in the indicator ”Conditional Accuracy”, Sun-Shine is slightly lower than the latest DeepSeek-R1.
-Also, as shown in Table XI, we also calculated the average of different metrics for different large models on different categories.
-Sun-Shine is either first or second in all categories and all indicators.
-Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance.
-Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine.
-Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
-| LLM | Response Rate | ||||
-| China | Social | Human | STEM | Other | |
-| Sun-Shine | 0.5507 | 0.5588 | 0.5377 | 0.5481 | 0.5498 |
-| GPT-3.5-turbo | 0.0687 | 0.0741 | 0.0748 | 0.0806 | 0.0801 |
-| GPT-4o | 0.4932 | 0.5133 | 0.5107 | 0.4915 | 0.4957 |
-| GPT-o1-mini | 0.2050 | 0.2340 | 0.2333 | 0.2231 | 0.2241 |
-| Llama-3.18B | 0.1907 | 0.1998 | 0.2056 | 0.2082 | 0.2104 |
-| Qwen-2.5 | 0.2563 | 0.2703 | 0.2736 | 0.2441 | 0.2419 |
-| DeepSeek-R1 | 0.4206 | 0.4069 | 0.3884 | 0.3580 | 0.3482 |
-| LLM | Conditional Accuracy | ||||
-| China | Social | Human | STEM | Other | |
-| Sun-Shine | 0.3208 | 0.3367 | 0.3477 | 0.3457 | 0.3489 |
-| GPT-3.5-turbo | 0.2489 | 0.2605 | 0.2656 | 0.2710 | 0.2731 |
-| GPT-4o | 0.2845 | 0.2871 | 0.2846 | 0.2746 | 0.2775 |
-| GPT-o1-mini | 0.2636 | 0.2700 | 0.2716 | 0.2742 | 0.2728 |
-| Llama-3.18B | 0.2510 | 0.2704 | 0.2662 | 0.2620 | 0.2597 |
-| Qwen-2.5 | 0.2656 | 0.2655 | 0.2710 | 0.2866 | 0.2900 |
-| DeepSeek-R1 | 0.3209 | 0.3482 | 0.3524 | 0.3584 | 0.3573 |
-| LLM | Accuracy | ||||
-| China | Social | Human | STEM | Other | |
-| Sun-Shine | 0.1888 | 0.1742 | 0.1683 | 0.1614 | 0.1618 |
-| GPT-3.5-turbo | 0.0182 | 0.0201 | 0.0205 | 0.0277 | 0.0265 |
-| GPT-4o | 0.1594 | 0.1676 | 0.1634 | 0.1515 | 0.1535 |
-| GPT-o1-mini | 0.0475 | 0.0584 | 0.0598 | 0.0603 | 0.0606 |
-| Llama-3.18B | 0.0656 | 0.0628 | 0.0723 | 0.0662 | 0.0663 |
-| Qwen-2.5 | 0.0691 | 0.0723 | 0.0744 | 0.0702 | 0.0709 |
-| DeepSeek-R1 | 0.1863 | 0.1819 | 0.1705 | 0.1562 | 0.1534 |
-Ti-Safetybench: as shown in Fig.
-11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods.
-Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA.
-GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
-As shown in Table XII, it presents a comprehensive comparison of leading LLMs on the Ti-Safetybench, evaluating their performance across key safety-related categories.
-Notably, Sun-Shine outperforms all baseline models under both DA and CCA tasks.
-In the DA setting, Sun-Shine achieves the highest exact match (EM) score (0.292) and leads in IA, OFF, and UB, while maintaining a significantly higher filter rate (59.20%) than GPT-4o (38.59%) and others.
-In the CCA setting, Sun-Shine demonstrates even stronger results, ranking first across nearly all categories, including EM (0.483), IA (0.455), mental health (MH, 0.507), and UB (0.466).
-These results highlight Sun-Shine’s superior ability to generate safe, accurate, and instruction-aligned responses, setting a new benchmark in safety-aware language modeling.
-| LLM | Category | ||||||||
-| Filter | Invaild | EM | IA | MH | OFF | PH | PP | UB | |
-| Sun-Shine | DA | 59.20% | 0.292 | 0.379 | 0.204 | 0.255 | 0.258 | 0.307 | 0.372 |
-| GPT-4o | 38.59% | 0.144 | 0.357 | 0.241 | 0.199 | 0.246 | 0.296 | 0.358 | |
-| Llama-3.18B | 74.75% | 0.105 | 0.088 | 0.101 | 0.097 | 0.094 | 0.099 | 0.109 | |
-| GPT-3.5-turbo | 73.42% | 0.085 | 0.073 | 0.087 | 0.086 | 0.097 | 0.106 | 0.123 | |
-| Qwen-2.5 | 49.50% | 0.199 | 0.118 | 0.164 | 0.198 | 0.150 | 0.159 | 0.371 | |
-| GPT-o1-mini | 71.20% | 0.109 | 0.070 | 0.095 | 0.111 | 0.097 | 0.095 | 0.170 | |
-| DeepSeek-R1 | 51.14% | 0.287 | 0.373 | 0.232 | 0.250 | 0.297 | 0.309 | 0.371 | |
-| Sun-Shine | CCA | 16.60% | 0.483 | 0.455 | 0.507 | 0.448 | 0.509 | 0.336 | 0.466 |
-| GPT-4o | 34.55% | 0.478 | 0.429 | 0.429 | 0.287 | 0.255 | 0.387 | 0.401 | |
-| Llama-3.18B | 69.72% | 0.130 | 0.101 | 0.123 | 0.124 | 0.119 | 0.115 | 0.126 | |
-| GPT-3.5-turbo | 67.75% | 0.108 | 0.082 | 0.103 | 0.125 | 0.114 | 0.118 | 0.160 | |
-| Qwen-2.5 | 16.61% | 0.448 | 0.413 | 0.427 | 0.289 | 0.259 | 0.385 | 0.403 | |
-| GPT-o1-mini | 61.69% | 0.254 | 0.099 | 0.134 | 0.163 | 0.135 | 0.117 | 0.231 | |
-| DeepSeek-R1 | 28.14% | 0.558 | 0.457 | 0.500 | 0.429 | 0.511 | 0.339 | 0.457 | |
-Overall: Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1.
-While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance.
-This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
-5.2 Sun-Shine for Tibetan Culture
-As shown in Fig.
-12, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese.
-This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records.
-The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
-As shown in Fig.
-13(a), Fig.
-13(b), and Fig.
-13(c), we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English.
-Sun-Shine generates outputs that are notably neutral, emotionally grounded, and culturally restrained across all three languages.
-The Tibetan response is respectful and sincere, avoiding ornate metaphors or religious references.
-In Chinese, the model articulates gratitude with clarity and warmth, emphasizing familial support without exaggeration.
-The English version maintains a reflective tone, expressing appreciation without invoking religious or hyperbolic imagery.
-This demonstrates Sun-Shine’s alignment with real-world communication norms and its ability to generalize across linguistic and cultural contexts.
-In contrast, GPT-4o exhibits a distinctly elevated and ceremonial tone.
-Its Tibetan and Chinese outputs are imbued with grandiloquent phrases and spiritual imagery.
-For example, the model refers to the father as a “beacon of happiness” and ”born on the edge of happiness,” which, while poetic, may not align with typical expressions of filial appreciation.
-The English version similarly includes metaphor-heavy language, such as “the victory banner is my prayer,” which resembles religious scripture more than natural expression.
-This suggests a stylistic bias toward exalted, sometimes deific, portrayals.
-DeepSeek-R1, on the other hand, appears to have an even stronger inclination toward religious and cultural iconography.
-It references “Mañjuśrī’s Father” and constructs its Tibetan output with vocabulary and structure closely mirroring classical Buddhist literature.
-The Chinese response is elegant yet heavily stylized, and the English output reads like a translated verse from scripture rather than a naturally composed poem.
-While aesthetically refined, this stylistic preference may reflect overfitting to religious texts in training data, limiting the model’s adaptability to general-purpose use.
-In summary, Sun-Shine’s responses are more balanced and less prone to embellishment, demonstrating cultural sensitivity and instruction alignment.
-GPT-4o and DeepSeek-R1, while fluent and creative, tend to overgenerate in stylistically loaded contexts, risking a mismatch with the intended pragmatic tone.
-This comparison underscores Sun-Shine’s strength in producing appropriate, grounded, and respectful language across multiple languages and cultural domains.
-5.3 Downstream Task
-Tibetan Language Proficiency Test: we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
-| LLM | A | B | C | D |
-| Sun-Shine | 63 | 61 | 44 | 42 |
-| GPT-4o | 45 | 47 | 53 | 51 |
-| DeepSeek-R1 | 59 | 56 | 67 | 66 |
-| Qwen-2.5 | 41 | 43 | 53 | 54 |
-As shown in Table XIII, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics.
-This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our benchmark, TIB-STC, which includes a richer collection of ancient Tibetan texts.
-Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan).
-BELU [40] and ROUGE [41] are used to evaluate this task.
-As shown in Table XIV, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower.
-Table 12 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
-| LLM | B-4 | R-1 | R-2 | R-l |
-| Sun-Shine | 42.16 | 65.78 | 48.91 | 51.58 |
-| GPT-4o | 38.85 | 61.12 | 44.33 | 47.02 |
-| DeepSeek-R1 | 41.85 | 65.49 | 47.73 | 51.04 |
-| Qwen-2.5 | 35.20 | 58.59 | 41.25 | 44.73 |
-6 Conclusion
-In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC benchmark, marking significant progress in Tibetan language processing.
-Sun-Shine, trained on the comprehensive TIB-STC benchmark, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios.
-TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data.
-Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan.
-Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our benchmark, TIB-STC, in providing a high-quality ancient Tibetan corpus.
-This work advances low-resource language models and underscores the importance of culturally specific benchmarks.
-Acknowledgment
-This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.